### v1

# Charge functions for high dimensional partitions

Abstract
To construct a BPS algebra with representations furnished by n-dimensional partitions, the first step is to construct the eigenvalue of the Cartan operators acting on them. The generating function of the eigenvalues are called the charge function. It has an important property that for each partition, the poles of the function correspond to the projection of the boxes which can be added to or removed from the partition legally. The charge functions of lower dimensional partitions, i.e., Young diagrams for 2D, plane partitions for 3D and solid partitions for 4D, are already given in the literature. In this paper, we propose an expression of the charge function for arbitrary odd dimensional partitions and have it proved for 5D case. Some explicit numerical tests for 7D and 9D case are also conducted to confirm our formula.
1 Introduction
Partitions, as fundamental combinatorial objects, have long played an important role in both mathematics and physics. The study of integer partitions – ways of writing a positive integer as a sum of positive integers – has deep connections to number theory, representation theory and statistical mechanics andrews1998theory . In 2D, Young diagrams provide a geometric visualization of partitions and arise naturally in the representation theory of symmetric groups and in the geometry of Hilbert schemes of points on surfaces nakajima1999lectures ; okounkov2007random . Their generating function is given by the celebrated Euler product formula, and they serve as building blocks for many integrable systems and random matrix models mehta2004random ; gross1980possible .
The natural generalization to 3D, known as plane partitions macmahon1898memoir , corresponding to stacking boxes in the positive octant of $\mathbb{Z}^{3}$ subject to non-increasing conditions along three axes. Plane partitions are intimately related to the geometry of $\mathbb{C}^{3}$, the simplest toric Calabi-Yau threefolds (CY3), where they count the equivalent Bogomol’nyi-Prasad-Sommerfield (BPS) states of D6-D0-branes in string theories ooguri2009crystal ; okounkov2006quantum ; iqbal2008quantum . The generating function of plane partitions is given by the MacMahon function, which also appears in the context of topological string theory and Donaldson-Thomas invariants donaldson1998gauge ; kontsevich2011cohomological ; maulik2006gromov ; maulik2019quantum .
For 4D, the analogous objects are solid partitions macmahon1912ix . They can be visualized by placing boxes in the positive corner of a 4D space. Physically, they enumerate equivariant BPS states of D8-D0-branes on the Calabi-Yau fourfolds (CY4), $\mathbb{C}^{4}$, generalizing the correspondence between plane partitions and D6-D0-states on $\mathbb{C}^{3}$ nekrasov2019magnificent ; bonelli2023adhm . Solid partitions are notoriously difficult to enumerate and their elusive generating function is believed to encode the partition function of D8-D0-branes on $\mathbb{C}^{4}$. The conjectured generating function of solid partitions introduced in macmahon2004combinatory failed at level 6, giving 141 instead of the true value 140 (see developments in nekrasov2020magnificent ; nekrasov2019magnificent ; bonelli2023adhm ; szabo2023instanton ; kimura2023double ; nekrasov2024global ; cao2018zero ).
A BPS algebra emerges by rendering the space of BPS states with an algebraic structure, as shown by Harvey and Moore harvey1996algebras ; harvey1998algebras . The BPS algebra has three types of operators: Cartan operators, creation operators and annihilation operators. They can be combined respectively into three generating fields of a complex spectral parameter $u$. For each partition, the eigenvalues of the Cartan generating operators are packaged into a meromorphic function of $u$, known as the charge function li2020quiver ; galakhov2024charging , which in turn governs the action of the creation and annihilation operators. With a certain ansatz on the actions of operators of the BPS algebra on the representation vector labeled by an $n$-dimensional partition, one may determine the BPS algebra with the help of its representation theory. This is introduced as a bootstrap procedure in li2020quiver . A key property of the charge function, which is essential for the BPS algebra, is that its poles are in one-to-one correspondence with the box positions where a box can be added by a creation operator or removed by an annihilation operator. So to construct a BPS algebra, we need to derive the appropriate charge function. If one has the charge function determined by this crucial property, and furthermore the action of creation/annihilation operators on the representation space, it is possible to write down the algebraic relation of the BPS algebra.
For plane partitions, this property forces the charge function to assume a factorized form: it is simply a product over all boxes with each contributing a basic rational function which is the core ingredient defining the BPS algebra. This algebra is the affine Yangian of $\mathfrak{gl}_{1}$ describing the BPS states of D6-D0-branes on $\mathbb{C}^{3}$ with the representation space labeled by plane partitions prochazka2016 ; rapcak2020cohomological ; galakhov2024charging . It is also known to be isomorphic to the central extension of Spherical degenerate double affine Hecke algebra (SHc) introduced by Shffmann and Vasserot in schiffmann2013cherednik ; arbesfeld2013presentation , to describe the equivalent cohomology of the instanton moduli space of $\mathcal{N}=2$ 4D gauge theories. This algebra precisely describe the algebraic structure behind Nekrasov instanton partition functions nekrasov2003seiberg with the Omega background and has been used to prove the 4D/2D correspondence proposed by Alday, Gaiotto and Tachikawa (AGT correspondence alday2010liouville ) and its various generalization kanno2013extended ; matsuo2014construction ; bourgine2016holomorphic ; bourgine2016coherent ; bourgine2019note .
Currently, the corresponding BPS algebra for solid partitions remains unknown, and in nekrasov2020magnificent this novel algebraic structure is termed as Mama algebra. In a significant advance, Galakhov and Li illuminatingly resolve the problem of the charge function in 4D case galakhov2024charging . [^2] A straightforward generalization of the simple form of charge functions in 3D to 4D can not work, since the charge function does not reproduce the correct poles structure associated with the addable and removable boxes for solid partitions. Some extra contributions from certain 4-box and 5-box clusters are introduced apart from the contributions from single boxes. They prove the constructed charge function satisfies all the required properties for any solid partition by checking all local pictures explicitly.
In the line of this development, we aim to advance this program of constructing charge functions for higher dimensional partitions, to explore the corresponding algebraic structure. Consequently, we conjecture an expression of the charge function for any odd dimensional partition and prove it for 5D case. We also conduct some explicit numerical tests for 7D and 9D to verify the formula. For higher even dimensional cases, we leave it for the future work.
This paper is organized as follows. In section 2, we review some necessary concepts about partitions in arbitrary dimension. In section 3, we conjecture an expression of the charge function for any odd dimension. Section 4 serves as a preparation for the following proof. In section 5, we demonstrate that the conjectured formula indeed satisfies the property of the charge function, as long as Lemma 5 holds. Then in section 6, we give a complete proof for 5D case. For higher odd dimensions, we offer a partial proof, and perform Monte Carlo sampling tests for 7D and 9D. We conclude in section 7. The details of the proof for Lemma 1-4 are collected in appendix A.
2 $n$-dimensional partitions
Let us show some basic concepts about partitions in this section following galakhov2024charging .
$\Delta^{(n)}$ is an $n$-dimensional partition (which is sometimes referred as molten crystal in physics), if it satisfies the melting rule: For any box $\vec{\square}\in\mathbb{Z}^{n}_{\geq 0}$, if there exists $\vec{\square}^{\prime}\in\Delta^{(n)}$ such that $\vec{\square}^{\prime}=\vec{\square}+\vec{e}_{k}$ for any $k=1,2,\cdots,n$, then $\vec{\square}\in\Delta^{(n)}$.
We denote the canonical basis in such a positive corner of $n$-dimensional space by
$\vec{e}_{i}=\left(\delta_{1i},\delta_{2i},\dots,\delta_{ni}\right),\quad i=1,2,\dots,n,$ (2.1)
where $\delta_{ij}$ is the Kronecker delta. We will use the notation:
$\vec{E}\equiv\sum_{i}^{n}\vec{e}_{i}$ (2.2)
in our later discussion.
For an $n$-dimensional partition $\Delta^{(n)}$, the Calabi-Yau condition reads:
$\sum_{i=1}^{n}h_{i}=0,$ (2.3)
where the complex number $h_{i}$ is called weight or flavor parameter. [^3] These parameters characterize the equivalent toric action on CYn as $(x_{1},x_{2},\cdots,x_{n})\mapsto(e^{h_{1}}x_{1},e^{h_{2}}x_{2},\cdots,e^{h_{n}}x_{n})$. There are two sets related to the partition $\Delta^{(n)}$ that are important
$\operatorname{A}_{\Delta^{(n)}}\subset\mathbb{Z}^{n}_{\geq 0}\quad\text{and}\quad\operatorname{R}_{\Delta^{(n)}}\subset\mathbb{Z}^{n}_{\geq 0},$ (2.4)
which respectively include all the boxes in $\mathbb{Z}^{n}_{\geq 0}$ that can be added to $\Delta^{(n)}$ or removed from $\Delta^{(n)}$ and the resulting partitions remain $n$-dimensional partitions. Let us introduce a projection action [^4] In this paper, we are accustomed to denote the coordinate component of a box $\vec{\square}$ by $l_{i}$.:
$c:\quad\vec{\square}=\sum_{i=1}^{n}l_{i}\vec{e}_{i}\longmapsto\sum_{i=1}^{n}l_{i}h_{i}.$ (2.5)
We assume that the weight parameter $h_{1,2,\cdots,n}$ are generic complex numbers that satisfy (2.3). The projection $c$ identifies points in $\mathbb{Z}^{n}_{\geq 0}$ that differ by multiples of the vector $\vec{E}$. So (2.5) it is not a unique decomposition of $c$. We can also define a unique component $l^{\prime}_{i}$ of c by letting the smallest component equals to zero.
$l^{\prime}_{i}=l_{i}-min\{l_{i},i={}1,2,..,n\},$ (2.6)
Then equivalently,
$c=\sum_{i=1}^{n}l^{\prime}_{i}h_{i},\quad l^{\prime}_{i}\geq 0.$ (2.7)
We will use the notation of $\ell$ and $\ell^{\prime}$ also as an action: $\ell(c)=\ell$, $\ell^{\prime}(c)=\ell^{\prime}$.
3 Conjecture: charge functions for odd dimensional partitions
After some careful consideration, we find that it is possible to construct the charge functions $\psi(u)$ for odd dimensional partitions $\Delta^{(n)}$. [^5] The even dimensional cases higher than 4D have more nuances, and we leave it for the future work. The charge function is supposed to satisfy the following properties: 1. $\psi_{\Delta^{(n)}}(u)$ is a meromorphic function of $u$ and only has simple poles. 2. All the poles of $\psi_{\Delta^{(n)}}(u)$ are in one-to-one correspondence with the projected vector $c$ of the boxes $\vec{\square}\in\mathrm{A}_{\Delta^{(n)}}\cup\mathrm{R}_{\Delta^{(n)}}$. Before presenting our formula, we first review the charge functions corresponding to lower dimensional partitions following galakhov2024charging .
Ordinary partitions
For 2D partitions (Young diagrams), the Calabi-Yau condition reads
$h_{1}+h_{2}=0.$ (3.1)
The charge function for the 2D partition $\Delta^{(2)}$ is :
$\psi_{\Delta^{(2)}}(u)=\frac{1}{u}\prod_{\vec{\square}\in\Delta^{(2)}}\varphi_{1}^{(2)}\left(u-c(\phi_{1})\right)\prod_{\phi_{2}\in\Delta^{(2)}}\varphi_{2}^{(2)}\left(u-c(\phi_{2})\right)\prod_{\phi_{3}\in\Delta^{(2)}}\varphi_{3}^{(2)}\left(u-c(\phi_{3})\right),$ (3.2)
where
$\varphi_{1}^{(2)}(u)=\frac{1}{(u-h_{1})(u-h_{2})},\quad\varphi_{2}^{(2)}(u)=u^{2},\quad\varphi_{3}^{(2)}(u)=\frac{1}{u^{2}}.$ (3.3)
and $\phi_{p}=\left\{\vec{\square},\vec{\square}+\vec{e}_{s_{1}},\vec{\square}+\vec{e}_{s_{2}},\cdots,\vec{\square}+\vec{e}_{s_{p}}\right\}$ is a $p$-box cluster, and the action of $c$ is defined as
$c(\phi_{p})=c(\vec{\square})+\sum_{i=1}^{p-1}h_{s_{i}}.$ (3.4)
Plane partitions
For 3D partitions (Plane partitions), the Calabi-Yau condition is
$h_{1}+h_{2}+h_{3}=0.$ (3.5)
The charge function for the plane partition $\Delta^{(3)}$ is :
$\psi_{\Delta^{(3)}}(u)=\frac{1}{u}\prod_{\vec{\square}\in\Delta^{(3)}}\varphi^{(3)}_{1}(u-c(\vec{\square})),$ (3.6)
where $\varphi_{1}^{(3)}(u)$ is the bonding factor:
$\varphi^{(3)}_{1}(u)=\prod_{i=1}^{3}\frac{u+h_{i}}{u-h_{i}}.$ (3.7)
For 3D case, the corresponding BPS algebra is the affine Yangian of $\mathfrak{gl}_{1}$ prochazka2016 .
Solid partitions
For 4D partitions (Solid partitions), the Calabi-Yau condition reads
$h_{1}+h_{2}+h_{3}+h_{4}=0.$ (3.8)
The charge function for the solid partition $\Delta^{(4)}$ is :
$\psi_{\Delta^{(4)}}(u)=\frac{1}{u}\prod_{\vec{\square}\in\Delta^{(4)}}\varphi_{1}^{(4)}\left(u-c(\phi_{1})\right)\prod_{\phi_{4}\in\Delta^{(4)}}\varphi_{4}^{(4)}\left(u-c(\phi_{4})\right)\prod_{\phi_{5}\in\Delta^{(4)}}\varphi_{5}^{(4)}\left(u-c(\phi_{5})\right),$ (3.9)
where
$\varphi_{1}^{(4)}(u)=\frac{\prod_{i=1}^{4}(u+h_{i})\prod_{1\leq i<j\leq 4}(u-h_{i}-h_{j})}{\prod_{i=1}^{4}(u-h_{i})},\quad{\varphi_{4}^{(4)}}(u)=\frac{1}{u^{2}},\quad{\varphi_{5}^{(4)}}(u)=u^{2}.$ (3.10)
and $c(\phi_{p})$ still take the form as (3.4). Equivalently,
$\varphi_{1}^{(4)}(u)=\frac{\prod_{1\leq i<j<k\leq 4}(u-h_{i}-h_{j}-h_{k})\prod_{1\leq i<j\leq 4}(u-h_{i}-h_{j})}{\prod_{i=1}^{4}(u-h_{i})}.$ (3.11)
Conjecture formula for arbitrary odd dimension:
We conjecture the form of the charge function in dimension $n=2K+1,K\geq 1,$
$\psi_{\Delta^{(n)}}(u)=\psi_{0}(u)\psi^{\prime}_{\Delta^{(n)}}(u),$ (3.12)
where
$\psi_{0}=\frac{1}{u},$ (3.13)
$\psi^{\prime}_{\Delta^{(n)}}(u)=\prod_{\vec{\square}\in\Delta^{(n)}}\varphi_{1}(u-c(\vec{\square}))\prod_{m=2}^{K}\prod_{\phi_{2m}\subset\Delta^{(n)}}\varphi_{2m}(u-c(\phi_{2m})).$ (3.14)
where $\phi_{2m}=\left\{\vec{\square},\vec{\square}+\vec{e}_{s_{1}},\vec{\square}+\vec{e}_{s_{2}},\cdots,\vec{\square}+\vec{e}_{s_{2m-1}}\right\}$ is a $2m$-box cluster, and $c(\phi_{p})$ still take the form as (3.4).
$c(\phi_{2m})=c(\vec{\square})+\sum_{i=1}^{2m-1}h_{s_{i}},$ (3.15)
and the exact form of the factor is as below:
$\varphi_{1}(u)=\frac{\prod_{m=1}^{K}\prod_{1\leq l_{1}<l_{2}<\cdots<l_{2m}\leq 2K+1}(u-\sum_{i=1}^{2m}h_{l_{i}})}{\prod_{i=1}^{2K+1}(u-h_{i})},$ (3.16)
$\varphi_{2m}(u)=\frac{1}{u}.$ (3.17)
It is easy to tell for $n=3$, i.e. $K=1$, (3.14) has no contribution from clusters, and (3.12) identifies with the known formula for plane partitions as shown in (3.6).
In the next section, we prove the formula proposed above indeed satisfies the properties of charge functions.
4 Definitions
Some useful sets
We define a set of all possible partitions in $n$ dimension, and a set of points in the projected space,
$P_{n}\coloneqq\left\{\Delta^{(n)}\ \text{for any number of boxes}\right\}\subset 2^{\mathbb{Z}_{\geq 0}^{n}},$ (4.1)
$\mathcal{P}\coloneqq\{c=\sum_{i=1}^{n}l_{i}h_{i}\bigm|l_{i}\in\mathbb{Z}\}\cong\mathbb{Z}^{n-1}.$ (4.2)
We define the union $C_{\Delta^{(n)}}\coloneqq\operatorname{A}_{\Delta^{(n)}}\cup\ \text{R}_{\Delta^{(n)}}$ including all the possible positions to add to or remove from a partition $\Delta^{(n)}$. We are interested in its projection space, so we define the set
$D_{\Delta^{(n)}}\coloneqq\left\{c=\sum_{i}l_{i}h_{i}\in\mathcal{P}\ \bigm|\ \sum_{i}l_{i}\vec{e}_{i}\in C(\Delta^{(n)})\right\}.$ (4.3)
Note that $\left|C(\Delta^{(n)})\right|=\left|D_{\Delta^{(n)}}\right|$. And we introduce another set to collect the simple pole of the partition
$\mathcal{SP}(\Delta^{(n)})\coloneqq\left\{c\ |\ c\text{ is a simple pole of }\psi_{\Delta^{(n)}}(u)\right\}.$ (4.4)
We define the set of admissible partitions at $\vec{\square}$, denoted by $G(\vec{\square})$, as the collection of all $n$-dimensional partitions for which there exists an addable or removable position that can be projected to $c(\vec{\square})$. That is,
$G(\vec{\square})=\{\Delta^{n}\in P_{n}\ |\ \exists\tilde{\vec{\square}}\in C(\Delta^{n}),\tilde{c}(\tilde{\vec{\square}})=c(\vec{\square})\}.$ (4.5)
Potential Function
We introduce the potential function to represent the order of the pole at point $c$,
$\omega_{0,\Delta^{(n)}}(c)=\begin{cases}m&\text{m-th order pole,}\\ 0&\text{no poles or zeros,}\\ -m&\text{m-th order zero.}\end{cases}$ (4.6)
The precise expression $\omega_{0,\Delta^{(n)}}(c)$ obtained from (3.14) is as follows. The base term is:
$\omega_{0,\Delta^{(n)}}(c)=\delta_{0,c}+\omega_{\Delta^{(n)}}(c).$ (4.7)
Given a set of box $S\subseteq\mathbb{Z}_{\geq 0}^{n}$, we further define:
[…]
where the individual components are defined as follows:
$\omega_{S,1}\coloneqq\sum_{\vec{\square}\in S}\left(\sum_{i=1}^{n}\delta_{\tilde{c},c(\vec{\square})+h_{i}}-\sum_{m=1}^{K}\,\sum_{0\leq l_{1}\leq\cdots\leq l_{2m}\leq n}\delta_{\tilde{c},c(\vec{\square})+\sum_{i=1}^{2m}h_{l_{i}}}\right),$ (4.9)
$\omega_{\phi_{2m}}\coloneqq\delta_{\tilde{c},c(\phi_{2m})},$ (4.10)
$\omega_{S,\phi_{2m}}\coloneqq\sum_{m=2}^{K}\,\sum_{\phi_{2m}\subseteq S}\omega_{\phi_{2m}}.$ (4.11)
We define the potential function for vector $\vec{\square}$:
$\omega_{S}(\vec{\square})\coloneqq\omega_{S}(c(\vec{\square})).$ (4.12)
Note that we have translation invariance for $\omega_{S}(\vec{\square})$, after translating every element in set $S$ and the specific box position $\vec{\square}$ along an arbitrary $n$-dimensional vector in $\mathbb{Z}^{n}$ to get $\tilde{S}$ and $\tilde{\vec{\square}}$ (with the requirement that all components of any element in the translated set $\tilde{S}$ are positive), for such a translation, we have:
$\omega_{\tilde{S}}(\tilde{\vec{\square}})=\omega_{S}(\vec{\square}).$ (4.13)
For disjoint sets $S_{1}$ and $S_{2}$ (i.e., $S_{1}\cap S_{2}=\emptyset$), the potential function satisfies additivity with cluster correction:
$\omega_{S_{1}\cup S_{2}}(c)=\omega_{S_{1}}(c)+\omega_{S_{2}}(c)+\omega_{cluster(S_{1},S_{2})}(c).$ (4.14)
Among them, the term $\omega_{cluster(S_{1},S_{2})}$ originates from the contributions of clusters set of $\mathcal{K}_{S_{1}\bowtie S_{2}}$ that satisfy the condition specified in this section: some boxes belong to set $S_{1}$ while the other part belongs to set $S_{2}$.
By decomposing the potential into two contributions from single boxes and clusters, we rewrite this as:
$\omega_{\text{cluster}(S_{1},S_{2})}=\left(\omega_{S_{1}\cup S_{2},1}-\omega_{S_{1},1}-\omega_{S_{2},1}\right)+\left(\omega_{S_{1}\cup S_{2},\phi_{2m}}-\omega_{S_{1},\phi_{2m}}-\omega_{S_{2},\phi_{2m}}\right).$ (4.15)
The first term vanishes identically, while the second term can be expressed as a sum over clusters crossing $S_{1}$ and $S_{2}$, defined as set of relevant clusters $\mathcal{K}_{S_{1}\bowtie S_{2}}$ :
$\omega_{S_{1}\cup S_{2},\phi_{2m}}-\omega_{S_{1},\phi_{2m}}-\omega_{S_{2},\phi_{2m}}=\sum_{\phi_{2m}\in\mathcal{K}_{S_{1}\bowtie S_{2}}}\omega_{\phi_{2m}}.$ (4.16)
Substituting this back to (4.15), we obtain the final expression of $\omega_{{cluster}}$:
$\omega_{{cluster}(S_{1},S_{2})}(\vec{\square})=\sum_{\phi_{2m}\in\mathcal{K}_{S_{1}\bowtie S_{2}}}\omega_{\phi_{2m}}(\vec{\square}).$ (4.17)
$d$-neighbor
A configuration $c$ is called a $d$-neighbor of $c^{\prime}$ (where $d<n$), denoted as $c^{\prime}\xrightarrow{d}c$ (we sometimes omit $d$ and write $c^{\prime}\leftrightarrow c$), if and only if:
$c=c^{\prime}+\sum_{i=1}^{d}h_{n_{i}},$ (4.18)
for some index set $\{n_{i}\}$ with $n_{i}\in\{1,2,\dots,n\}$. It is straightforward to show that the neighborhood relation is symmetric in the sense of dimension complement:
$c^{\prime}\xrightarrow{d}c\iff c\xrightarrow{n-d}c^{\prime}.$ (4.19)
Note that we have properties:
  1. 1.
    Given $\vec{\square}$:
    $\exists S,\,\omega_{S}(c^{\prime})\neq\omega_{S-\vec{\square}}(c^{\prime})\implies c(\vec{\square})\leftrightarrow c^{\prime},$ (4.20)
  2. 2.
    The neighborhood relation is equivalent to the following conditions:
    $c\leftrightarrow c^{\prime}\iff\forall i,\,l_{i}^{\prime}(c-c^{\prime})\in\{0,1\}.$ (4.21)
    which is further equivalent to:
    $\forall i,j,\,l_{i}(c-c^{\prime})-l_{j}(c-c^{\prime})\in\{-1,0,1\}.$ (4.22)
Bisect operation
Now we define an operation that divides the partition set $\Delta^{(n)}$ resides into two subsets, $L{(\Delta^{(n)},\vec{\square})}$ and $\Delta^{(n)}-L{(\Delta^{(n)},\vec{\square})}$.
For a partition $\Delta^{(n)}$, define $L$ as operation that yields a collection of boxes $L$, where each box satisfies the condition that all its components are greater than or equal to the corresponding components of box $\vec{\square}=(l_{1},l_{2},...,l_{i})$:
$L{(\Delta^{(n)},\vec{\square})}\coloneqq\Delta^{(n)}\cap\{\tilde{\vec{\square}}=\sum_{i=1}^{n}\tilde{l_{i}}\vec{{e_{i}}}\,|\,\tilde{l_{i}}-l_{i}\geq 0,\,\forall\tilde{l_{i}}\}$. (4.23)
Hypercube
A $d$-dimensional hypercube $HC$ with its origin $\vec{\square}$ is defined as a subspace with $2^{d}$ elements in $\mathbb{Z}^{n}$:
$HC^{(d)}(\vec{\square},\{e_{n_{i}}\}_{i=1}^{d})\coloneqq\{\tilde{\vec{\square}}\,|\,\tilde{\vec{\square}}=\vec{\square}+\sum_{i=1}^{d}\delta_{i}\vec{e}_{n_{i}},\delta_{i}=0,1\}.$ (4.24)
(a) Schematic diagram of the definition of bisect operation and hypercube.
(b) Schematic diagram of the definition of surface set.
Figure 1: Schematic diagram of definition in 2D.
Surface set
We now define the set of positions located on the surface (which means it is the position of the smallest components among those unoccupied position with the same c) of the partition. For a fixed dimension $d\leq n$ and a given set of basis vectors $\{e_{n_{i}}\}$ (where $i=1,\dots,d$, $n_{i}\in[1,n]$), we define the set of $d$-dimensional surface points as
$F(d,\{\vec{e}_{n_{i}}\})\coloneqq\left\{\vec{\square}\,\middle|\,\begin{aligned} &\vec{\square}=\sum_{i=1}^{n}l_{i}\vec{e}_{i}\notin\Delta^{(n)};\\ &\,l_{n_{j}}\neq 0\ (\forall j\in[1,d]),\ l_{k}=0\ (\forall k\notin{\{n_{j}\}}),\quad(d<n)\\ &\vec{\square}-\vec{E}\in\Delta^{(n)}\quad(d=n)\end{aligned}\right\}.$ (4.25)
That is to say, for $\vec{\square}\in F(d,\{\vec{e}_{n_{i}}\})$, we mean $\vec{\square}$ is a position on the surface of partition with ndn-d components equal to zero and the set of nonzero directions is $\{\vec{e}_{n_{i}}\}$.
Taking the union over all possible dimensions and all index sets, given $\Delta^{(n)}$ we define the complete set of surface positions:
$F\coloneqq\bigcup_{d=0}^{n}\bigcup_{\{n_{i}\}}F(d,\{\vec{e}_{n_{i}}\}).$ (4.26)
Note that there exists a natural bijection between the boxes (positions) in $F$ and the elements of the projection space $\mathcal{P}$; thus, the set $F$ precisely captures the surface structure of interest.
Conjecture
Using the notation introduced above, we can rewrite our conjecture as: 1. $\psi_{\Delta^{(n)}}(u)$ has only simple poles. 2. $D_{\Delta^{(n)}}=\mathcal{SP}_{\Delta^{(n)}}$, which is equivalent to $\Delta^{(n)}\in G(\vec{\square})\,\,\Longleftrightarrow\,\,\omega_{0,\Delta^{(n)}}(\vec{\square})=1.$
5 Proof of the Conjecture
In this section, we start with introducing some lemmas, and prove the conjecture basing on these lemmas.
5.1 Lemmas
Lemma 1.
(proved in Appendix A.1) For $\forall$ $\Delta^{(n)}$ and $\vec{\square}$, after bisect operation $L$, the remaining boxes in $\Delta^{(n)}-L$ still form a partition.
$\Delta^{(n)}-L(\Delta^{(n)},\vec{\square})\in P_{n}\,.$ (5.1)
Lemma 2.
(proved in Appendix A.2) For $\forall$ $\Delta^{(n)}$ and $\vec{\square}\in F(d,\{\vec{e}_{n_{i}}\})$, The potential at $\vec{\square}$ of original partition $\Delta^{(n)}$ is equal to that in processed partition $\Delta^{(n)}-\tilde{L}+\widetilde{HC}$}.
[…]
Where for later convenience, we denote.
$\widetilde{HC}\equiv HC^{(d)}\left(\vec{\square}-\sum_{i=1}^{d}\vec{e}_{n_{i}},\{\vec{e}_{n_{i}}\}\right)\cap\Delta^{(n)},\vec{\square}\in F(d,\{\vec{e}_{n_{i}}\}),$ (5.3)
$\widetilde{L}\equiv L(\Delta^{(n)},\vec{\square}-\vec{E}).$ (5.4)
In the definition here, we specify an $n$-dimensional vector $\vec{\square}$ as the position for performing the bisect operation $\tilde{L}$ and generating the hypercube $\widetilde{HC}$.
Lemma 3.
(proved in Appendix A.3) For $\forall$ $\Delta^{(n)}$ and $\vec{\square}\in F(d,\{\vec{e}_{n_{i}}\})$. The potential contribute by the clusters in $\tilde{\mathcal{K}}\equiv\mathcal{K}_{\widetilde{HC}\bowtie(\Delta^{(n)}-\tilde{L})}$ is zero.
$\omega_{cluster(\widetilde{HC},\Delta^{(n)}-\tilde{L})}(\vec{\square})=0.$ (5.5)
Lemma 4.
(proved in Appendix A.4) For $\forall$ $\Delta^{(n)}$ and $\vec{\square}\in F(d,\{\vec{e}_{n_{i}}\})$, $d<n$.
$\omega_{\Delta^{(n)}-\tilde{L}}(\vec{\square})=0$ (5.6)
Lemma 5.
(proved for $n=5$ and discussed for higher n in section 6) For $\forall$ $\Delta^{(n)}\subseteq HC^{(d)}(\vec{0},\{\vec{e}_{n_{i}}\})$.
$1,$ $\quad\text{if }\Delta^{(n)}\in G\left(\sum_{i=1}^{d}\vec{e}_{n_{i}}\right),$ (5a)
$\leq 0,$ $\quad\text{if }\Delta^{(n)}\notin G\left(\sum_{i=1}^{d}\vec{e}_{n_{i}}\right).$ (5b)
5.2 Main proof
We proceed by mathematical induction. The conjecture is immediate in $|\Delta^{(n)}|=0$. Assume that our conjecture holds for all cases where $|\Delta^{(n)}|<N$. Now, consider the case where $|\Delta^{(n)}|=N$. Given a vector $\vec{\square}\in F(d,\{\vec{e}_{n_{i}}\})$, $\vec{\square}=\sum_{i=1}^{n}l_{i}\vec{e_{i}}$, we analyze the potential function $\omega_{\Delta^{(n)}}(\vec{\square})$ in two scenarios based on the dimension $d$.
We begin by writing the potential function with Lemma 2:
[…]
The third term is zero according to Lemma 3, then
[…]
First, consider the case where $d=n$, noted that:
$\tilde{\vec{\square}}\equiv\vec{\square}-\vec{E}\in A_{\Delta^{(n)}-\tilde{L}}$, (5.9)
since $\tilde{\vec{\square}}+\vec{e}_{i}\in\tilde{L}$, $\tilde{\vec{\square}}-\vec{e}_{i}\in\Delta^{(n)}-\tilde{L}$ and $\tilde{\vec{\square}}\notin\Delta^{(n)}-\tilde{L}$. We also have $|\Delta^{(n)}-\tilde{L}|<N$ because at least $\tilde{\vec{\square}}$ is removed. And we know from Lemma 1 that $\Delta^{(n)}-\tilde{L}$ is a partition. Our conjecture at lower level guarantees that $\omega_{\Delta^{(n)}-\tilde{L}}(\vec{\square})=\omega_{\Delta^{(n)}-\tilde{L}}(\tilde{\vec{\square}})=1$. (5.8) is now:
$\omega_{\Delta^{(n)}}(\vec{\square})=1+\omega_{\widetilde{HC}}(\vec{\square})=1+\omega_{\widetilde{HC}-[\vec{\square}-\vec{E}]}(\vec{E})=\omega_{0,\widetilde{HC}-[\vec{\square}-\vec{E}]}(\vec{E}).$ (5.10)
The second equal holds because translate invariance see (4.13), $\widetilde{HC}-[\vec{\square}-\vec{E}]$ means that the new hypercube obtained by translating the hypercube along $\vec{\square}-\vec{E}$ and its origin exactly located at $\vec{0}$.
Second, consider the case where $d<n$. The first term in (5.8) equals zero By Lemma 4 :
$\omega_{\Delta^{(n)}}(\vec{\square})=\omega_{\widetilde{HC}}(\vec{\square})=\omega_{\widetilde{HC}-[\vec{\square}-\sum_{i=1}^{d}\vec{e}_{n_{i}}]}(\sum_{i=1}^{d}\vec{e}_{n_{i}})=\omega_{0,\widetilde{HC}-[\vec{\square}-\sum_{i=1}^{d}\vec{e}_{n_{i}}]}(\sum_{i=1}^{d}\vec{e}_{n_{i}}).$ (5.11)
Which has the same form as (5.10) Finally, by Lemma 5,
$\omega_{0,\widetilde{HC}-[\vec{\square}-\sum_{i=1}^{d}\vec{e}_{n_{i}}]}(\sum_{i=1}^{d}\vec{e}_{n_{i}})=1\iff\widetilde{HC}-[\vec{\square}-\sum_{i=1}^{d}\vec{e}_{n_{i}}]\in G(\sum_{i=1}^{d}\vec{e}_{n_{i}}).$ (5.12)
On the other hand, for any $\vec{\square}\in F(d,\{\vec{e}_{i}\})$, $\Delta^{(n)}\in G(\vec{\square})\iff\widetilde{HC}-[\vec{\square}+\sum_{i=1}^{d}\vec{e}_{n_{i}}]\in G(\sum_{i=1}^{d}\vec{e}_{n_{i}})$ because the melting rule only involves those boxes in $\widetilde{HC}$. As a result, we have the equivalence $\omega_{\Delta^{(n)}}(\vec{\square})=1\iff\Delta^{(n)}\in G(\vec{\square})$. Since we can see that $\omega_{\Delta^{(n)}}(\vec{\square})\leq 1$ from Lemma 5, this proves the conjecture.
6 Discussion of Lemma 5
For convenience, we repeat the statement of Lemma 5:
Lemma 5
$\forall$ $\Delta^{(n)}\subseteq HC^{(d)}(\vec{0},\{\vec{e}_{n_{i}}\})$,
$\omega_{0,\Delta^{(n)}}(\sum_{i=1}^{d}\vec{e}_{n_{i}})=\begin{cases}1&\Delta^{(n)}\in G(\sum_{i=1}^{d}\vec{e}_{n_{i}}),\\ \leq 0&\Delta^{(n)}\notin G(\sum_{i=1}^{d}\vec{e}_{n_{i}}).\end{cases}$ (6.1)
Figure 2: All six unique partitions of the hypercube ${HC}^{(2)}$ in 5D.
To study Lemma 5, We write $HC^{(d)}\left(\vec{0},\{\vec{e}_{n_{i}}\}_{i=1}^{d}\right)=HC^{(d)}$ for short. We first claim that: if $\Delta^{(n)}\subseteq HC^{(d)}$,
$\Delta^{(n)}\in G(\sum_{i=1}^{d}\vec{e}_{n_{i}})\iff\begin{cases}\{\emptyset,\{\vec{0}\},HC^{(n)}-\{\vec{E}\},HC^{(n)}\}&d=n,\\ \{\{HC^{(d)}-{\sum_{i=1}^{d}\vec{e}_{n_{i}}}\},HC^{(d)}\}&d\leq n.\end{cases}$ (6.2)
It is easy to check by the melting rule. As an example, Fig.2 shows the projections of all six unique partitions of the hypercube with $d=2$ in the 5D case ${HC}^{(2)}$. It can be observed that the pole order at the target is 1 only for Partition 5 and Partition 6 in Fig.2, which is consistent with the conclusion that they belong to $G(\sum_{i=1}^{d}\vec{e}_{n_{i}})$.
6.1 Some analytically check of Lemma 5
First we give a proof of Lemma 5a:
$\Delta^{(n)}\in G(\sum_{i=1}^{d}\vec{e}_{n_{i}})\Rightarrow\omega_{0,\Delta^{(n)}}(\sum_{i=1}^{d}\vec{e}_{n_{i}})=1.$ (6.3)
First, it is straightforward to show:
$w_{0,\{0\}}(\vec{E})=w_{0,\emptyset}(\vec{E})=1,$ (6.4)
$w_{0,HC^{(d)}}=w_{0,HC^{(d)}-\{\sum_{i=1}^{d}\vec{e}_{n_{i}}\}}.$ (6.5)
We can also prove
$w_{0,HC^{(d)}}=1.$ (6.6)
The number of nmn-m neighbor of $c(\sum_{i=1}^{d}e_{n_{i}})=\sum_{i=1}^{d}h_{n_{i}}$ is:
$|\{c\ |\ c\xrightarrow{m}\sum_{i=1}^{d}h_{n_{i}},c\in HC^{(d)}\}|=C_{d}^{m}.$ (6.7)
Noting that $\forall\,k\in\{3,5\dots 2k-1\}$, there’s one cluster centered at each nkn-k neighbor of $\sum_{i=1}^{d}e_{n_{i}}$ that contributes a pole to $\sum_{i=1}^{d}e_{n_{i}}$. In a word, each $n-m,(1\leq m\leq n-1)$ neighbor contributes $(-1)^{n+1}$ poles. This is the same predicted by (7.1).
Case 1: $n\neq d$
$w_{HC^{(d)}}=\sum_{m=1}^{d}C_{d}^{m}(-1)^{m+1}=1,$ (6.8)
$w_{0,HC^{(d)}}=w_{HC^{(d)}}.$ (6.9)
Case 2: $n=d$
$w_{HC^{(n)}}=\sum_{m=1}^{n-1}C_{d}^{m}(-1)^{m+1}=0,$ (6.10)
$w_{0,HC^{(n)}}=1+w_{HC^{(n)}}=1.$ (6.11)
Thus we have complete the proof of Lemma 5a.
6.2 Numerical proof of Lemma 5
For the d-dimensional hypercubes satisfying the requirements in Lemma 5, we numerically enumerated all unique partitions for dd=1 to 5, proving that Lemma 5 holds rigorously for all 5D case (totally $3+6+20+168+7581$ unique cases).
For the 7 and 9 dimensional case, Because the Dedekind number M(7) & M(9) is too large ($2.4147\times 10^{12}$ and a 42-digit value calculated in 2023 D9 ), we performed Monte Carlo sampling, verifying partitions with different numbers of boxes for dd=1 to $n$. For all existing sampling results, the upper bound predictions of Lemma 5 for potential fully meet the requirements.
Fig.3(a)&3(b) shows the distribution of the number of unique partitions as a function of the number of boxes composing the partitions, for ${HC}^{(4)}$ and ${HC}^{(5)}$ respectively. The maximum number of unique partitions occurs when the number of boxes is $2^{d-1}$.
Fig.3(c)&3(d) show the order of the pole at the target position corresponding to partitions with different numbers of constituent boxes in ${HC}^{(4)}$ and ${HC}^{(5)}$. The size of the bubble represents the quantity of unique partitions with particular number of boxes and target pole order. It can be observed that for $d=4$, the pole order is 1 only for 15 and 16 boxes case (the fully occupied case and the case with one missing box), while for $d=5$, the pole order is 1 when the number of boxes is 0/1/31/32 in $d=5$ case.
(a) ${HC}^{(4)}$ partition distribution
(b) ${HC}^{(5)}$ partition distribution
(c) ${HC}^{(4)}$ pole order at target position
(d) ${HC}^{(5)}$ pole order at target position
Figure 3: Numerical results for $n=5$. Upper panel: Distributions of unique partition counts vs. box numbers for ${HC}^{(4)}$ and ${HC}^{(5)}$, respectively. Lower panel: Target pole orders for ${HC}^{(4)}$ and ${HC}^{(5)}$ partitions with different box numbers. Bubble size denotes the count of unique partitions for each (box number, pole order) pair. For $d=4$, pole order=1 only for 15/16 boxes (fully occupied/one missing box); for $d=5$, pole order=1 for 0/1/31/32 boxes.
(a) Monte Carlo sampling results for
$d=n=7$
(b) Monte Carlo sampling results for
$d=n=9$
Figure 4: Monte Carlo sampling results for 7D and 9D case, where the sample results are in good agreement with the description of Lemma 5.
Figure 5: Visualization of a special $n=5$ partition (200 boxes). Each subplot shows its projection onto $h_{1},h_{2},h_{3}$, with horizontal rightward as increasing $h_{4}$ and vertical downward as increasing $h_{5}$. Red/green squares denote the positions which can add new boxes or remove existing boxes (melting rule), and black dots represent simple poles of the charge function. Their perfect coincidence confirms consistent judgments between the two methods, verifying the universality of our method for general cases.
Fig.4(a)&4(b) shows the Monte Carlo sampling results for $n=d=7$ and $n=d=9$ case, it can be clearly seen that the results for the samples are in good agreement with the description of Lemma 5, providing numerical confidence for our proof. Fig.5 presents the visualization of a special $n=5$ partition (comprising 200 boxes). Each subplot shows its projection onto the first three dimensions ($h_{1},h_{2},h_{3}$), where the horizontal rightward direction represents the increasing order of $h_{4}$, and the vertical downward direction represents the increasing order of $h_{5}$. In the figure, the red and green squares denote the positions which can add new boxes or remove existing boxes via the melting rule, while the black dots represent the simple poles of the charge function. It can be observed that the two perfectly coincide, indicating consistent judgment results between the two methods and verifying the universality of our method for general cases.
7 Summary and discussion
In this paper, we achieve a breakthrough by successfully constructing the charge function (3.12) that is universally applicable to any odd-dimensional partitions. A critical foundation of our proof lies in Lemma 5, whose validity is indispensable for ensuring the rigor and correctness of the entire theoretical framework. Only when Lemma 5 holds can the charge function effectively fulfill its designed role. To consolidate this foundational result, we not only prove for Lemma 5a, but also conduct comprehensive, numerical validations to corroborate the reliability of Lemma 5a and Lemma 5b.
We perform an exhaustive enumeration of all unique partitions in $d$-dimensional hypercubes for dimensions $d=1$ to $5$ in $\mathbb{Z}_{\geq 0}^{5}$. This exhaustive search confirms that Lemma 5 holds for all 5D cases (encompassing a total of 7778 unique cases), thus finishing the proof for 5D case. For higher-dimensional cases, exhaustive enumeration becomes computationally intractable due to the exponential growth of the partition space ($2^{2^{n}}$). Instead, we adopt a Monte Carlo sampling approach, which systematically verifies partitions with varying numbers of boxes across the dimensional range from $d=1$ to the target dimension $n$ (i.e., $7$ and $9$), ensuring broad coverage of possible partitions.
Notably, all numerical results, whether from exhaustive enumeration ($n=5$) or Monte Carlo sampling ($n=7,9$), consistently demonstrate that the upper bound predictions for the potential derived from Lemma 5 fully meet the required theoretical conditions. Building on this validated foundation of Lemma 5, we further rigorously prove that the constructed charge function satisfies the crucial properties 1 and 2, which are essential for accurately capturing the correct pole structure of the system. Collectively, our theoretical construction of the charge function, Lemma 5a, and extensive numerical verifications (covering $d=1$ to $5$ via exhaustive enumeration and $d=7,9$ via Monte Carlo sampling) provide compelling evidence for the validity and robustness of our proposed framework.
However, it is not easy to generalize our result (3.14) to even dimensional cases for the following two reasons.
1. Odd-order product terms induce asymmetric distribution
(3.14) can be formally written in an approximation form,
$\psi^{\prime}_{\Delta^{(n)}}(u)\sim\prod_{\vec{\square}\in\Delta^{(n)}}\varphi(u-c(\vec{\square})),$ (7.1)
where
$\varphi(u)=\frac{\prod_{m=1}^{K}\prod_{1\leq l_{1}<l_{2}<\cdots<l_{2m}\leq 2K+1}(u-\sum_{i=1}^{2m}h_{l_{i}})}{\prod_{m=1}^{K}\prod_{1\leq l_{1}<l_{2}<\cdots<l_{2m-1}\leq 2K+1}(u-\sum_{i=1}^{2m-1}h_{l_{i}})}.$ (7.2)
For a large partition, (7.1) differs from (3.14) only in the contribution from clusters at the surface of the partition. There are odd number of integers $s\in[1,n-1]$ if n is even, which implies that terms of the form
$\prod\limits_{1\leq l_{1}<l_{2}<\cdots<l_{s}\leq n}\left(u-\sum_{i=1}^{s}h_{l_{i}}\right),$ (7.3)
cannot be evenly distributed between the numerator and denominator in (7.1).
2. Pole contribution breaks sign symmetry for even $n$
The pole contribution from $\vec{\square}$ must be +1+1 to a 1-neighbor of $c(\vec{\square})$, and 1-1 to an (n1)(n-1)-neighbor. For even $n$, this implies the pole contribution from $\vec{\square}$ to a $d$-neighbor of $c(\vec{\square})$ cannot follow the conjectured pattern $(-1)^{d}$.
In spite of the even dimensional issue mentioned above, which we leave for future work, our result serves as a foundational step toward constructing BPS algebras for Calabi-Yau $n$-folds. Just as the charge function for $n=3$ leads to the affine Yangian of $\mathfrak{gl}_{1}$ and the $n=4$ case motivates the Solid Algebra, our formula provides the necessary eigenvalue data to bootstrap the algebra generators for $n=5$ and beyond.
Acknowledgements. The authors thank Yutaka Matsuo, Jean-Emile Bourgine, Rui-Dong Zhu, Zifan Chen and Martijn Kool for helpful discussions. K.Z. (Hong Zhang) is supported by a classified fund of Shanghai city.
Appendix A Proof of Lemma 1-4
A.1 Lemma 1
For $\forall$ $\Delta^{(n)}$ and $\vec{\square}$, after bisect operation $L$, the remaining boxes in $\Delta^{(n)}-L$ still form a partition.
$\Delta^{(n)}-L(\Delta^{(n)},\vec{\square})\in P_{n}\,.$ (A.1)
Proof.
For $\forall\ \tilde{\vec{\square}}\in\Delta^{(n)}-L(\Delta^{(n)},\vec{\square})$ and $i\in\{1,2,...,n\}$, the melting rule in $\Delta^{(n)}$ implies:
$\tilde{\vec{\square}}-\vec{e}_{i}\in\Delta^{(n)},$ (A.2)
We also have:
$\tilde{\vec{\square}}-\vec{e}_{i}\notin L(\Delta^{(n)},\vec{\square}),$ (A.3)
due to the definition of $L$ (4.23). Thus we have
$\tilde{\vec{\square}}-\vec{e}_{i}\in\Delta^{(n)}-L(\Delta^{(n)},\vec{\square}),$ (A.4)
which is the melting rule of $\Delta^{(n)}-L(\Delta^{(n)},\vec{\square})$ that proof the Lemma by definition. ∎
A.2 Lemma 2
For $\forall$ $\Delta^{(n)}$ and $\vec{\square}\in F(d,\{\vec{e}_{n_{i}}\})$, The potential at $\vec{\square}$ of original partition $\Delta^{(n)}$ is equal to that in processed partition $\Delta^{(n)}-\tilde{L}+\widetilde{HC}$}.
[…]
Proof.
First, we start with the inclusion relation that the hypercube $\widetilde{HC}$ is a subset of $\widetilde{L}$, i.e.,
$\widetilde{HC}\subset\widetilde{L}$. (A.6)
Based on this inclusion, we can decompose $\widetilde{L}$ into the disjoint union of $\widetilde{L}-\widetilde{HC}$ and $\widetilde{HC}$, which gives
$\widetilde{L}=(\widetilde{L}-\widetilde{HC})\cup\widetilde{HC}$. (A.7)
Next, consider an arbitrary vector $\tilde{\vec{\square}}\in\widetilde{L}-\widetilde{HC}$. For all indices $i$, the component-wise condition holds due to $\tilde{\vec{\square}}\in\tilde{L}$:
$\widetilde{l}_{i}-(l_{i}-1)\geq 0.$ (A.8)
Since $\vec{\square}\notin\widetilde{HC}$, there exists at least one index $i$ such that the component difference satisfies
$\widetilde{l}_{i}-(l_{i}-1)\geq 2\Leftrightarrow\widetilde{l}_{i}-l_{i}\geq 1.$ (A.9)
On the other hand, because $\vec{\square}\notin\Delta^{(n)}$ and using the melting rule, there exists some index $j$ where
$\widetilde{l}_{j}-l_{j}\leq-1.$ (A.10)
Combining these two results, we find that there exist indices $i,j$ such that the difference of component differences is bounded below by 2:
$(\widetilde{l}_{i}-l_{i})-(\widetilde{l}_{j}-l_{j})\geq 2.$ (A.11)
(4.22) then implies that the n1n-1 vector $c(\vec{\square})$ is not a neighbor of c(~)c($\tilde{\vec{\square}}$), denoted as
[…]
Now, take any intermediate set $I$ satisfying $\Delta^{(n)}-\widetilde{L}+\widetilde{HC}\subsetneq I\subset\Delta^{(n)}$. By the neighborhood non-equivalence established above and (4.20), the potential function $w_{I}(\vec{\square})$ remains unchanged when removing any $\tilde{\vec{\square}}\in\widetilde{L}-\widetilde{HC}$, i.e.,
$w_{I}(\vec{\square})=w_{I-\tilde{\vec{\square}}}(\vec{\square})\quad\forall\ \tilde{\vec{\square}}\in\widetilde{L}-\widetilde{HC}$. (A.13)
By iteratively removing all elements of $\widetilde{L}-\widetilde{HC}$ from $I$ and using the invariance of the potential function, we finally obtain the desired equality:
[…]
A.3 Lemma 3
For $\forall$ $\Delta^{(n)}$ and $\vec{\square}\in F(d,\{\vec{e}_{n_{i}}\})$. The potential contribute by the clusters in $\tilde{\mathcal{K}}=\mathcal{K}_{\widetilde{HC}\bowtie(\Delta^{(n)}-\tilde{L})}$ is zero,
$\omega_{cluster(\widetilde{HC},\Delta^{(n)}-\tilde{L})}(\vec{\square})=0.$ (A.15)
Proof.
Suppose $\phi_{2m}$ is a cluster contributing to
[…]
We first prove the following key claim:
$\vec{\square}_{c}\in\Delta^{(n)}-\tilde{L}\quad\text{and}\quad\exists!k,\,\vec{\square}_{c}+\vec{e}_{k}\in\tilde{L}$. (A.17)
[…]
From above, we immediately derive the component-wise relation for the vector difference:
$l_{i}(\vec{\square}-\vec{\square}_{c})=\begin{cases}1&i=k,\\ \leq 0&i\neq k,\end{cases}$ (A.18)
[…]
Using this definition, we expand the cluster potential function step-by-step:
[…]
$=\sum_{\phi_{2m}\in\tilde{\mathcal{K}}}\delta_{c(\vec{\square})-c(\vec{\square}_{c}),\sum_{i=1}^{2m-1}hs_{i}}\quad(\text{rearranging the delta function})$
note that the first term under $\delta$ has 1 largest component while the second has 2m12m-1 ($m>2$), so they can’t be equal
$\omega_{\text{cluster}}(\vec{\square})=0.$ (A.20)
A.4 Lemma 4
For $\forall$ $\Delta^{(n)}$ and $\vec{\square}\in F(d,\{\vec{e}_{n_{i}}\})$, $d<n$,
$\omega_{\Delta^{(n)}-\tilde{L}}(\vec{\square})=0.$ (A.21)
Proof.
We begin by leveraging the condition $d<n$ (where $d$ denotes the dimension of the surface set containing $\vec{\square}$). By the definition of $d$-dimensional surface points (see Definition of surface set 4), this dimension condition implies there exists at least one index $i$ such that the $i$-th component of $\vec{\square}$ is zero, i.e.,
$d<n\Rightarrow\forall j\neq n_{i},\quad l_{j}(\vec{\square})=0.$ (A.22)
Next, consider an arbitrary vector $\tilde{\vec{\square}}\in\Delta^{(n)}-\widetilde{L}$ (the remaining partition after removing $\widetilde{L}$ via the bisect operation). By the definition of L (4.23), there exists an integer k, such that:
$l_{n_{k}}(\tilde{\vec{\square}})-l_{n_{k}}\left(\vec{\square}-\sum e_{n_{i}}\right)<-1.$ (A.23)
We now analyze the component-wise difference $l_{j}(\tilde{\vec{\square}})-l_{j}(\vec{\square})$:
Substituting $\vec{\square}=\vec{\square}-\sum e_{n_{i}}+\sum e_{n_{i}}$, we find:
$l_{n_{k}}(\tilde{\vec{\square}})-l_{n_{k}}(\vec{\square})=l_{n_{k}}(\tilde{\vec{\square}})-\left(l_{n_{k}}\left(\vec{\square}-\sum e_{n_{i}}\right)+1\right)<-1$ (A.24)
Combining this with the fact that $l_{j}(\vec{\square})=0$, we further derive the difference between the $j$-th and $n_{k}$-th components of $\tilde{\vec{\square}}-\vec{\square}$:
$l_{j}(\tilde{\vec{\square}}-\vec{\square})-l_{n_{k}}(\tilde{\vec{\square}}-\vec{\square})>1+l_{j}(\tilde{\vec{\square}})>1.$ (A.25)
By the property of the neighborhood relation (4.22), this implies the n1n-1 vector C(~)C($\tilde{\vec{\square}}$) and $C(\vec{\square})$ are not neighbors, denoted as:
$C(\tilde{\vec{\square}})\nleftrightarrow C(\vec{\square}).$ (A.26)
The property of neighbors (4.20) gives a key consequence for the potential function: for any subset $I\subseteq\Delta^{(n)}-\widetilde{L}$ and any $\vec{\square}^{\prime}\in I$, removing $\vec{\square}^{\prime}$ from $I$ does not change the potential at $\vec{\square}$.
$\forall I\subseteq\Delta^{(n)}-\widetilde{L},\quad\vec{\square}^{\prime}\in I,\quad w_{I-\vec{\square}^{\prime}}(\vec{\square})=w_{I}(\vec{\square}).$ (A.27)
We can iteratively apply this result by removing all elements from $\Delta^{(n)}-\widetilde{L}$ one by one. Eventually, we reduce $I$ to the empty set $\emptyset$, and since the potential function of the empty set at any position is zero ($w_{\emptyset}(\vec{\square})=0$), we conclude:
$w_{\Delta^{(n)}-\widetilde{L}}(\vec{\square})=w_{\emptyset}(\vec{\square})=0.$ (A.28)

### v2

@@
-Charge functions for high dimensional partitions
+Charge functions for odd dimensional partitions
 Abstract
 To construct a BPS algebra with representations furnished by n-dimensional partitions, the first step is to construct the eigenvalue of the Cartan operators acting on them.
-The generating function of the eigenvalues are called the charge function.
+The generating function of the eigenvalues is called the charge function.
 It has an important property that for each partition, the poles of the function correspond to the projection of the boxes which can be added to or removed from the partition legally.
@@
 We conclude in section 7.
-The details of the proof for Lemma 1-4 are collected in appendix A.
+The details of the proofs for Lemma 1-4 are collected in appendix A.
 2 $n$-dimensional partitions
@@
 We are interested in its projection space, so we define the set
-$D_{\Delta^{(n)}}\coloneqq\left\{c=\sum_{i}l_{i}h_{i}\in\mathcal{P}\ \bigm|\ \sum_{i}l_{i}\vec{e}_{i}\in C(\Delta^{(n)})\right\}.$ (4.3)
-Note that $\left|C(\Delta^{(n)})\right|=\left|D_{\Delta^{(n)}}\right|$. And we introduce another set to collect the simple pole of the partition
+$D_{\Delta^{(n)}}\coloneqq\left\{c=\sum_{i}l_{i}h_{i}\in\mathcal{P}\ \bigm|\ \sum_{i}l_{i}\vec{e}_{i}\in C_{\Delta^{(n)}}\right\}.$ (4.3)
+Note that $\left|C_{\Delta^{(n)}}\right|=\left|D_{\Delta^{(n)}}\right|$. And we introduce another set to collect the simple pole of the partition
 $\mathcal{SP}(\Delta^{(n)})\coloneqq\left\{c\ |\ c\text{ is a simple pole of }\psi_{\Delta^{(n)}}(u)\right\}.$ (4.4)
@@
 That is,
-$G(\vec{\square})=\{\Delta^{n}\in P_{n}\ |\ \exists\tilde{\vec{\square}}\in C(\Delta^{n}),\tilde{c}(\tilde{\vec{\square}})=c(\vec{\square})\}.$ (4.5)
+$G(\vec{\square})=\{\Delta^{(n)}\in P_{n}\ |\ \exists\,\tilde{\vec{\square}}\in C_{\Delta^{(n)}},\tilde{c}(\tilde{\vec{\square}})=c(\vec{\square})\}.$ (4.5)
 Potential Function
@@
 For a partition $\Delta^{(n)}$, define $L$ as operation that yields a collection of boxes $L$, where each box satisfies the condition that all its components are greater than or equal to the corresponding components of box $\vec{\square}=(l_{1},l_{2},...,l_{i})$:
-$L{(\Delta^{(n)},\vec{\square})}\coloneqq\Delta^{(n)}\cap\{\tilde{\vec{\square}}=\sum_{i=1}^{n}\tilde{l_{i}}\vec{{e_{i}}}\,|\,\tilde{l_{i}}-l_{i}\geq 0,\,\forall\tilde{l_{i}}\}$. (4.23)
+$L{(\Delta^{(n)},\vec{\square})}\coloneqq\Delta^{(n)}\cap\{\tilde{\vec{\square}}=\sum_{i=1}^{n}\tilde{l_{i}}\vec{{e_{i}}}\,|\,\tilde{l_{i}}-l_{i}\geq 0,\,\forall\,i\}.$ (4.23)
 Hypercube
-A $d$-dimensional hypercube $HC$ with its origin $\vec{\square}$ is defined as a subspace with $2^{d}$ elements in $\mathbb{Z}^{n}$:
+A $d$-dimensional hypercube $HC$ with its origin $\vec{\square}$ is defined as a subspace with $2^{d}$ elements in $\mathbb{Z}_{\geq 0}^{n}$:
 $HC^{(d)}(\vec{\square},\{e_{n_{i}}\}_{i=1}^{d})\coloneqq\{\tilde{\vec{\square}}\,|\,\tilde{\vec{\square}}=\vec{\square}+\sum_{i=1}^{d}\delta_{i}\vec{e}_{n_{i}},\delta_{i}=0,1\}.$ (4.24)
@@
 6.1 Some analytically check of Lemma 5
-First we give a proof of Lemma 5a:
+We give a proof of Lemma 5a:
 $\Delta^{(n)}\in G(\sum_{i=1}^{d}\vec{e}_{n_{i}})\Rightarrow\omega_{0,\Delta^{(n)}}(\sum_{i=1}^{d}\vec{e}_{n_{i}})=1.$ (6.3)
 First, it is straightforward to show:
-$w_{0,\{0\}}(\vec{E})=w_{0,\emptyset}(\vec{E})=1,$ (6.4)
+$w_{0,\{\vec{0}\}}(\vec{E})=w_{0,\emptyset}(\vec{E})=1,$ (6.4)
 $w_{0,HC^{(d)}}=w_{0,HC^{(d)}-\{\sum_{i=1}^{d}\vec{e}_{n_{i}}\}}.$ (6.5)
@@
 $|\{c\ |\ c\xrightarrow{m}\sum_{i=1}^{d}h_{n_{i}},c\in HC^{(d)}\}|=C_{d}^{m}.$ (6.7)
-Noting that $\forall\,k\in\{3,5\dots 2k-1\}$, there’s one cluster centered at each nkn-k neighbor of $\sum_{i=1}^{d}e_{n_{i}}$ that contributes a pole to $\sum_{i=1}^{d}e_{n_{i}}$.
+Noting that $\forall\,k\in\{3,5\dots 2K-1\}$, there’s one cluster centered at each nkn-k neighbor of $\sum_{i=1}^{d}\vec{e}_{n_{i}}$ that contributes a pole to $\sum_{i=1}^{d}\vec{e}_{n_{i}}$.
 In a word, each $n-m,(1\leq m\leq n-1)$ neighbor contributes $(-1)^{n+1}$ poles.
-This is the same predicted by (7.1).
 Case 1: $n\neq d$
@@
 Figure 5: Visualization of a special $n=5$ partition (200 boxes).
-Each subplot shows its projection onto $h_{1},h_{2},h_{3}$, with horizontal rightward as increasing $h_{4}$ and vertical downward as increasing $h_{5}$.
+Each subplot shows its projection onto $\vec{e}_{1},\vec{e}_{2},\vec{e}_{3}$, with horizontal rightward as increasing $\vec{e}_{4}$ and vertical downward as increasing $\vec{e}_{5}$.
 Red/green squares denote the positions which can add new boxes or remove existing boxes (melting rule), and black dots represent simple poles of the charge function.
@@
 (Hong Zhang) is supported by a classified fund of Shanghai city.
-Appendix A Proof of Lemma 1-4
+Appendix A Proofs of Lemma 1-4
 A.1 Lemma 1
@@
 $\tilde{\vec{\square}}-\vec{e}_{i}\in\Delta^{(n)}-L(\Delta^{(n)},\vec{\square}),$ (A.4)
-which is the melting rule of $\Delta^{(n)}-L(\Delta^{(n)},\vec{\square})$ that proof the Lemma by definition. ∎
+which is equivalent to the melting rule of $\Delta^{(n)}-L(\Delta^{(n)},\vec{\square})$, thus proving the Lemma by definition. ∎
 A.2 Lemma 2